4/x+4/2x=-7
Solve for x

Answer by colliefan(242)

4/x+4/2x=-7
Get a common denominator.
8/2x+4/2x=-7
(8+4)/2x=-7
12/2x=-7
6x=-7
6x*1/6=-7*1/6
x*6*1/6=-7/6
x*1=-7/6
x=-7/6